the electronic potential in a region is v= 5x^2+3y^2+xy where V is volt. What is the electric field at the point (1,3). Express your answer in unit vector form
Answer #1

En a lon + 3x0+ os = 1076 + ty Va sx²+ - 3y = my E V dal au da av by + ay E flou+y) + +(6740)] P(1,3) (10x1+3). 16 x3 +1) Eye